OpenStudy (anonymous):

solve for the indicated letter f=2g,for g g=?

OpenStudy (anonymous):

g=?

OpenStudy (anonymous):

g=f/2

OpenStudy (anonymous):

g=f/2

OpenStudy (anonymous):

F=2G G=F/2 Answer. :)